OpenStudy (anonymous):

Please help me solve this problem! Suppose that 25 acres of land are purchased, but because of roads and wetland concerns, 6 2/3 acres cannot be developed into lots. how many 5/6 acre lots can the rest of the land be dividend in to

OpenStudy (anonymous):

Answer is 22. Change 6 2/3 and 25 to fraction form, you got 20/3 and 75/3. 75/3 - 20/3 = 55/3, which is the acres that can be developed. Now divide 55/3 by 5/6. \[\frac{ 55 }{ 3 } \div \frac{ 5 }{ 6} = \frac{ 55 }{ 3 } * \frac{ 6 }{ 5 }\] Cancels and you have 22 acres.

OpenStudy (anonymous):

how did you get the 75/3 @study100

OpenStudy (anonymous):

You want 20/3 and 25 to have the same denominator to subtract them, so think of it in terms of this: 25 * 1 = 25 correct? 3/3 = 1 25 * (3/3) = 75 /3 Now if you get the calculator and type 75/3 it would equate to 25.

OpenStudy (anonymous):

GREAT thank you!